The expected real rate of interest is 0.5%, actual inflation over the last year was -0.05%, and the nominal interest rate is currently 0.28%. According to the Fisher equation, what is the expected inflation (in %) over the next year, dPe? Round to 0.01%. E.g., if your answer is 3.145%, record it as 3.15
Answer #1

Expected Real Rate of Interest, r = 0.50%
Nominal Interest Rate, n = 0.28%

According to Fisher equation:

(1 + n) = (1 + r) * (1 + dPe)
(1 + dPe) = (1 + n) / (1 + r)
dPe = (n - r) / (1 + r)

dPe = (0.0028 - 0.0050) / (1 + 0.0050)
dPe = -0.0022 / 1.0050
dPe = -0.0022 or -0.22%

So, expected inflation over the next year is -0.22%